OpenStudy (anonymous):

if a^2+b^2=170 and ab=77 fine (a+b)^2. please teach me a quick method on how to solve this. thanks!

OpenStudy (anonymous):

as we know that (a+b)^2 = (a^2) + (b^2) + 2ab now it is given that a^2 + b^2 = 170 and ab = 77 (a+b)^2 = 170 + 2(77) = 170 + 154 = 324

OpenStudy (anonymous):

got it phoebeco

OpenStudy (anonymous):

well thanks for awarding me with a medal

OpenStudy (amistre64):

there are no "quick" methods; there are sensible methods at best

OpenStudy (amistre64):

seeing that they give you 2 equations with the same variables in them we can equate them to find out the as and bs that they have in common

OpenStudy (amistre64):

ab=77; we can solve this for either a or b a = 77/b or b = 77/a pick one of these options and use it as a value to replace in the other equation; ill use a = 77/b a^2 + b^2 = 170 (77/b)^2 + b^2 = 170 ; now we have 1 variable to worry about and can try to solve for b

OpenStudy (amistre64):

5929/ b^2 + b^2 = 170 ; *b^2 5929 + b^4 = 170b^2 b^4 -170b^2+5929 = 0 ; this is a quadratic in disguise, to see it let x = b^2 x^2 -170x+5929 = 0; solve for x and equate those solutions to x = b^2

OpenStudy (amistre64):

i get x=49 and x=121 from that... sooo x = b^2 49 = b^2 49 = (+-7)^2 x = b^2 121 = b^2 121 = (+-11)^2 try out all these options to weed out the extra ones

OpenStudy (amistre64):

a * b = 77 -7 * -11 = 77 a^2 + b^2 = 170 49 + 121 = 170 ; that works out a * b = 77 7 * 11 = 77 a^2 + b^2 = 170 49 + 121 = 170 ; that works out as well; :)

OpenStudy (amistre64):

(7+11)^2 18^2 18 18 --- 144 18 ---- 324 .... same conclusion :)

OpenStudy (anonymous):

that was a perfect answer amistre 64